2. A compound containin ng uranium and fluorine weighed 1.760 grams and was treated with excess calcium nitrate. The precipitate of CaF, which formed weighed 1.170 grams. What was the % of fluorine in the original sample?
Answer #1

The weight of CaF2 = 1.17 gm

Moles of CaF2 = 1.17 / 78.07 =  0.01498 Moles

Moles of Florine presents = 0.0299 Moles

Weight of Fluorine =  0.0299 x 18.99 = 0.5694 gm

% of Fluorine in the sample = 0.5694 x 100 /1.76 = 32.35 %